OpenStudy (anonymous):

Hey, whats the slope intercept again I totally forgot, is it the same as the Y-intercept? @iGreen

OpenStudy (igreen):

Slope intercept?

OpenStudy (igreen):

I think you made a typo.

OpenStudy (anonymous):

Well, here is the question.. What is the slope-intercept form of this equation? 3x + 9y = 18 @iGreen

OpenStudy (igreen):

Slope intercept form is \(y = mx + b\).

OpenStudy (anonymous):

OKay, so M is 3 and B is 9?

OpenStudy (igreen):

Just switch the values around. \(3x + 9y = 18\) Subtract 3x to both sides: \(9y = -3x + 18\) Divide 9 to all terms: \(y = \dfrac{-3}{9} + 2\) Simplify the fraction: \(y = -\dfrac{1}{3}+2\)

OpenStudy (anonymous):

Thank you, again.

OpenStudy (igreen):

No, for here you have to change the form to y = mx + b. Otherwise, you can't find the slope and y-intercept.

OpenStudy (igreen):

No problem.